Year to date, company O had earned a -2.10 % return. During the same time period, company v earned 8.00 % and company m earned 6.35%. If you have a portfolio made up of 40% company o, 30% company v, and 30% company m, what is your portfolio return?
Answer by Theo(13342)

company O earned -2.10% return.
company V earned 8% return.
company M earned 6.35% return.

your portfolio contains 40% of company O, 30% of company V, and 30% of company M.

what is the overall composite return of the portfolio?

let x equal the total amount invested in the portfolio.

the return on company O is equal to .4 * x * -2.10%.
the return on company V is equal to .3 * x * 8%.
the return on company M is equal to .3 * x * 6.35%.

the overall composite return would be .4 * x * -2.10% + .3 * x * 8% + .3 * x * 6.35% = 3.465% * x.

to test this out, let x be equal to any random number.

i chose 1000 dollars.

.4 * 1000 = 400
.3 * 1000 = 300
.3 * 1000 = 300

return on 400 = -2.10% = 400 * -2.10 / 100 = -8.4.
return on first 300 = 300 * 8 / 100 = 24.
return on second 300 = 300 * 6.35% = 300 * 6.35 / 100 = 19.05

total return = -8.4 + 24 + 19.05 = 34.65.
34.65 / 1000 = .03465 * 100 = 3.465%.

solution is confirmed to be good.

solution is that the overall composite portfolio return is 3.465%.